OpenStudy (anonymous):

kindly integrate this please. int. dx / x sqr (2ax - x^2) let x + 2a sin^2x

OpenStudy (anonymous):

\[\int\limits dx / x \sqrt{2ax - x^2}\] let x = 2ax sin^x

OpenStudy (anonymous):

let x = 2a sin^2x **

myininaya (myininaya):

is this right problem? \[\int\limits_{}^{}\frac{dx}{x \sqrt{2ax-x^2}} =\int\limits_{}^{}\frac{dx}{x \sqrt{-(x^2-2ax)}} \] \[\int\limits_{}^{}\frac{dx}{x \sqrt{(- (x^2- 2ax+a^2)+a^2}} =\int\limits_{}^{}\frac{dx}{x \sqrt{-(x-a)^2+a^2}}\]

myininaya (myininaya):



myininaya (myininaya):